Quantitative Aptitude Practice Question and Answer
8Q: If a trader sold two cars each at Rs. 325475 and gains 12% on the first and loses 12% on the second, then his profit or loss percent on the whole is ? 4012 05b5cc6ebe4d2b4197774ee78
5b5cc6ebe4d2b4197774ee78- 11.44% losstrue
- 214.4% lossfalse
- 30%false
- 414.4% profitfalse
- Show AnswerHide Answer
- Workspace
- SingleChoice
Answer : 1. "1.44% loss"
Explanation :
Answer: A) 1.44% loss Explanation: SP of each car is Rs. 325475, he gains 12% on first car and losses 12% on second car.In this case, there will be loss and percentage of loss is given by = [(profit%)(loss%)]/100 = (12)(12)/100 % = 1.44%
Q: In a race of 1000 m, A can beat by 100 m, in a race of 800m, B can beat C by 100m. By how many meters will A beat C in a race of 600 m ? 3962 05b5cc6e5e4d2b4197774ed42
5b5cc6e5e4d2b4197774ed42- 1127.5 mtrue
- 2254 mfalse
- 3184 mfalse
- 4212 mfalse
- Show AnswerHide Answer
- Workspace
- SingleChoice
Answer : 1. "127.5 m"
Explanation :
Answer: A) 127.5 m Explanation: When A runs 1000 m, B runs 900 m and when B runs 800 m, C runs 700 m. When B runs 900 m, distance that C runs = (900 x 700)/800 = 6300/8 = 787.5 m. In a race of 1000 m, A beats C by (1000 - 787.5) = 212.5 m to C. In a race of 600 m, the number of meters by which A beats C = (600 x 212.5)/1000 = 127.5 m.
Q: A bus is running at 60 kmph without any stoppage. But, with stoppages its average speed is 50 kmph. Find the time of stoppage by bus per hour. 3945 05d15f74644a3a75f13296b7d
5d15f74644a3a75f13296b7d- 115 minutes/hourfalse
- 220 minutes/hourfalse
- 312 minutes/hourfalse
- 410 minutes/hourtrue
- Show AnswerHide Answer
- Workspace
- SingleChoice
Answer : 4. "10 minutes/hour"
Q: A and B invests Rs.10000 each, A investing for 8 months and B investing for all the 12 months in the year. If the total profit at the end of the year is Rs.25000, find their shares? 3944 05b5cc7bbe4d2b419777509cd
5b5cc7bbe4d2b419777509cd- 110000 and 15000true
- 215000 and 10000false
- 35000 and 20000false
- 420000 and 5000false
- Show AnswerHide Answer
- Workspace
- SingleChoice
Answer : 1. "10000 and 15000"
Explanation :
Answer: A) 10000 and 15000 Explanation: As both A and B invest the same amounts, the ratio of their profits at the end of the year is equal to the ratio of the time periods for which they have invested. Thus, the required ratio of their profits = A : B = 8 : 12 = 2 : 3. Hence, share of A in the total profit = 2 x 25000/5 = Rs.10000 Similarly, share of B in the total profit = 3 x 25000/5 = Rs.15000
Q: In a regular week, there are 5 working days and for each day, the working hours are 8. A man gets Rs. 2.40 per hour for regular work and Rs. 3.20 per hours for overtime. If he earns Rs. 432 in 4 weeks, then how many hours does he work for ? 3935 05b5cc74fe4d2b4197774fb35
5b5cc74fe4d2b4197774fb35- 1145false
- 2165false
- 3175true
- 4135false
- Show AnswerHide Answer
- Workspace
- SingleChoice
Answer : 3. "175"
Explanation :
Answer: C) 175 Explanation: Given that, working days = 5 working hours = 8 A man get rupess per hour is Rs.2.40 So in one day the man get total rupees is 2.40 x 8 = 19.2 So in 5 days week the man get total rupees is 19.2 x 5 = 96 So in 4 week the man get total rupees is 96 x 4 = 384 So the man worked for = 160hours in 4 weeks But given that the man earned Rs.432 Hence remaning money is (432-384 = 48)which is earn by doing overtime work Overtime hours = 48/3.20 = 15 So total worked hours is = 15 + 160 = 175.
Q: What is the next number in the given Number Series? 5, 24, 94, 279, ? 3919 05b5cc6b9e4d2b4197774d73e
5b5cc6b9e4d2b4197774d73e- 1587false
- 2554true
- 3489false
- 4499false
- Show AnswerHide Answer
- Workspace
- SingleChoice
Answer : 2. "554"
Explanation :
Answer: B) 554 Explanation: The given number series follows a pattern that 5 x 5 – 1 = 24 24 x 4 – 2 = 94 94 x 3 – 3 = 279 279 x 2 – 4 = 554
Q: A man spend Rs. 810 in buying trouser at Rs. 70 each and shirt at Rs. 30 each. What will be the ratio of trouser and shirt when the maximum number of trouser is purchased ? 3916 05b5cc731e4d2b4197774f7b5
5b5cc731e4d2b4197774f7b5- 11:3false
- 22:1false
- 33:2true
- 42:3false
- Show AnswerHide Answer
- Workspace
- SingleChoice
Answer : 3. "3:2"
Explanation :
Answer: C) 3:2 Explanation: Let us assume S as number of shirts and T as number of trousersGiven that each trouser cost = Rs.70 and that of shirt = Rs.30Therefore, 70 T + 30 S = 810=> 7T + 3S = 81......(1)T = ( 81 - 3S )/7 We need to find the least value of S which will make (81 - 3S) divisible by 7 to get maximum value of TSimplifying by taking 3 as common factor i.e, 3(27-S) / 7In the above equation least value of S as 6 so that 27- 3S becomes divisible by 7 Hence T = (81-3xS)/7 = (81-3x6)/7 = 63/7 = 9 Hence for S, put T in eq(1), we getS = 81-7(9)/3 = 81-63 / 3 = 18/3 = 6.The ratio of T:S = 9:6 = 3:2.
Q: 8 couples (husband and wife) attend a dance show "Nach Baliye' in a popular TV channel ; A lucky draw in which 4 persons picked up for a prize is held, then the probability that there is atleast one couple will be selected is : 3902 05b5cc7d0e4d2b41977751218
5b5cc7d0e4d2b41977751218- 18/39false
- 215/39true
- 312/13false
- 4None of thesefalse
- Show AnswerHide Answer
- Workspace
- SingleChoice
Answer : 2. "15/39"
Explanation :
Answer: B) 15/39 Explanation: P( selecting atleast one couple) = 1 - P(selecting none of the couples for the prize) = 1-16C1× 14C1×12C1×10C116C4=1539

